# Diffusion approximation of the stationary distribution of a two-level single server queue

### v6

@@ 4.2 Preliminary lemmas
$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}+o(r),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r)+o(r),\quad i=1,2,$ (4.35)
$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+o(r),$ (4.36)
$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+o(r).$ (4.37)
@@
For each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$,
$e^{-\eta^{(r)}_{i}(r\theta_{i})\widehat{R}^{(r)}_{e}}=1-r\lambda_{i}\theta_{i}\widehat{R}^{(r)}_{e}-2^{-1}r^{2}\lambda_{i}^{3}\sigma_{e,i}^{2}r\theta_{i}^{2}(\widehat{R}^{(r)}_{e})^{2}+o(r^{2}),$ (4.38)
$e^{-\zeta^{(r)}(r\theta_{i})\widehat{R}^{(r)}_{s}}=1+\mu r\theta_{i}\widehat{R}^{(r)}_{s}-2^{-1}r^{i}\mu^{3}\sigma_{s}^{2}\theta_{i}^{2}(\widehat{R}^{(r)}_{s})^{2}+o(r^{2}),\qquad i=1,2,$ (4.39)
@@
$\hskip 34.44434pt=r(\lambda_{2}\theta_{2}-\lambda_{1}\theta_{1})\widehat{R}^{(r)}_{e}+2^{-1}r^{2}(\lambda_{2}^{3}\sigma_{e,2}^{2}r\theta_{2}^{2}-\lambda_{1}^{3}\sigma_{e,1}^{2}r\theta_{1}^{2})(\widehat{R}^{(r)}_{e})^{2}+o(r^{2}).$ (4.40)
$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s}}$
$\hskip 34.44434pt=\mu r(\theta_{2}-\theta_{1})\widehat{R}^{(r)}_{s}-2^{-1}r^{2}\mu^{3}\sigma_{s}^{2}(\theta_{2}^{2}-\theta_{1}^{2})(\widehat{R}^{(r)}_{s})^{2})+o(r^{2}),$ (4.41)
These lemmas are proved in Appendix A.4. From BAR (4.3) in Lemma 4.3, we aim to determine the limiting measures $\widetilde{\nu}_{1}$ and $\widetilde{\nu}_{2}$. This will be done step by step.
(1st Step)   We expand (4.3) using Lemmas 4.7 and 4.8. For this, we need the asymptotic behaviors concerning $\widehat{R}^{(r)}_{e}$ and $\widehat{R}^{(r)}_{s}$ under $\mathbb{P}^{(r)}_{s}$ and $\mathbb{P}^{(r)}_{e}$, respectively. To see them, we first put $\theta_{2}=\theta_{1}=\theta$ in (4.3), then
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o(r^{2})\big{)}\psi^{(r)}_{i}(\theta)$
$\quad=c_{1}\left(\mu r\theta+\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o(r^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta),$ (4.42)
where applying (4.35), (4.8) and Corollary 4.3 to (4.2) yields
$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+\theta r+o(r))\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r(\lambda_{2}-\lambda_{1})\theta(1+\theta r)\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r^{2})$
$\quad=\alpha^{(r)}_{e}e^{\theta\ell_{0}}r(\lambda_{2}-\lambda_{1})\theta\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]+O(r^{2}).$
Hence, dividing both sides of (4.3) by $r$ and letting $r\downarrow 0$, we have
$\lim_{r\downarrow 0}\Big{(}c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\Big{)}=0.$
Since $e^{r\theta\ell^{(r)}_{0}}=e^{(\ell_{0}+o(1))\theta}$ can be arbitrarily small by choosing small $\theta<0$, we have
$\lim_{r\downarrow 0}\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=0.$
and therefore, by Corollary 4.2, we have
$\lim_{r\downarrow 0}\mathbb{P}[L^{(r)}=0]=0.$ (4.43)
We next put $\theta_{1}=0$ and $\theta_{2}=\theta$ in (4.3). In this case,
$\frac{1}{2}c_{2}\sigma_{2}^{2}\big{(}(\beta_{2}-\theta)\mu\theta r^{2}+o(r^{2})\big{)}\psi^{(r)}_{2}(\theta)+D^{(r)}(0,\theta)=0,$
and, from (4.2),
@@
$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
@@
$=\alpha^{(r)}_{e}(e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}])$
$\quad+\alpha^{(r)}_{e}(\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]-e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})])$
$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}(\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})])$
$=\alpha^{(r)}_{e}r\theta e^{r\theta\ell^{(r)}_{0}}\big{(}\mu\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\lambda_{2}(1+r\theta)\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{)}+O(r^{2}),$
where the 3rd equation is obtained by (4.3), and the last one is by Corollary 4.3 and Lemma 4.8. Since $D^{(r)}(0,\theta)=O(r^{2})$ from (4.3), we have
$\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=O(r),\quad\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r),$ (4.44)
because $\mathbb{E}^{(r)}_{e}[\widehat{R}^{(r)}_{s}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$ and $\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ are nonnegative. These results lead to the following asymptotic results, which strengthens Corollary 4.3.
@@
$\lim_{r\downarrow 0}\mathbb{P}^{(r)}_{e}[L^{(r)}(0)=\ell^{(r)}_{0}]=0,\quad\lim_{r\downarrow 0}\mathbb{P}^{(r)}_{s}[L^{(r)}(0-)=\ell^{(r)}_{0}]=0,$ (4.45)
$\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{e}[(\widehat{R}^{(r)}_{s})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,\quad\lim_{r\downarrow 0}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0-)=\ell^{(r)}_{0})]=0.$ (4.46)

### v7

@@ 4.3 Expansion of the asymptotic BAR
$e^{r\theta_{i}\ell^{(r)}_{0}}=e^{\theta_{i}\ell_{0}}(1+o(\theta_{i}r)),\quad e^{r\theta_{i}(\ell^{(r)}_{0}+1)}=e^{\theta_{i}\ell_{0}}(1+\theta_{i}r+o(\theta_{i}r)),\quad i=1,2,$ (4.37)
$e^{r\theta_{1}\ell^{(r)}_{0}}-e^{r\theta_{2}\ell^{(r)}_{0}}=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)),$ (4.38)
$e^{r\theta_{1}(\ell^{(r)}_{0}+1)}-e^{r\theta_{2}(\ell^{(r)}_{0}+1)}$
$\quad=e^{\theta_{1}\ell_{0}}-e^{\theta_{2}\ell_{0}}+r(e^{\theta_{1}\ell_{0}}\theta_{1}-e^{\theta_{2}\ell_{0}}\theta_{2})+e^{(\theta_{1}\vee\theta_{2})\ell_{0}}(o(\theta_{1}r)+o(\theta_{2}r)).$ (4.39)
@@
Let $S_{n}(x)=\sum_{k=1}^{n}x^{k}$ for $x\in\mathbb{R}$ and integer $n\geq 1$, then, for each fixed $\theta_{1}$ and $\theta_{2}$, as $r\downarrow 0$, for $i=1,2$,
$e^{-\eta^{(r)}_{i}(r\theta_{i})\widehat{R}^{(r)}_{e}}=1-r\theta_{i}(\lambda^{(r)}_{i}\widehat{R}^{(r)}_{e})-2^{-1}(r\theta_{i})^{2}[(\lambda^{(r)}_{i}\sigma^{(r)}_{e,i})^{2}(\lambda^{(r)}_{i}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{i}\widehat{R}^{(r)}_{e})^{2}]$
$\qquad+S_{3}(\widehat{R}^{(r)}_{e})o((r\theta)^{2}),$ (4.40)
$e^{-\zeta^{(r)}(r\theta_{i})\widehat{R}^{(r)}_{s-}}=1+r\theta_{i}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-2^{-1}(r\theta_{i})^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]$
$\qquad+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}),$ (4.41)
@@
$\quad=r(\theta_{2}\lambda^{(r)}_{2}-\theta_{1}\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}+2^{-1}r^{2}[\theta_{2}^{2}(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}\lambda^{(r)}_{2}-\theta_{1}^{2}(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}\lambda^{(r)}_{1}]\widehat{R}^{(r)}_{e}$
$\qquad-2^{-1}r^{2}[\theta_{2}^{2}(\lambda^{(r)}_{2})^{2}-\theta_{1}^{2}(\lambda^{(r)}_{1})^{2}](\widehat{R}^{(r)}_{e})^{2}+S_{3}(\widehat{R}^{(r)}_{e})o((r\theta)^{2}),$ (4.42)
$e^{-\zeta^{(r)}(r\theta_{2})\widehat{R}^{(r)}_{s-}}-e^{-\zeta^{(r)}(r\theta_{1})\widehat{R}^{(r)}_{s-}}$
$\quad=r(\theta_{2}-\theta_{1})(\mu^{(r)}\widehat{R}^{(r)}_{s-})-2^{-1}r^{2}(\theta_{2}^{2}-\theta_{1}^{2})[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]$
$\qquad+S_{3}(\widehat{R}^{(r)}_{s-})o((r\theta)^{2}).$ (4.43)
These lemmas are proved in Appendix A.4.
We first compute $D^{(r)}(\theta,0)$.
$D^{(r)}(\theta,0)$ $=\alpha^{(r)}_{e}(1-e^{r\theta\ell^{(r)}_{0}})\mathbb{P}^{(r)}_{e}[L^{(r)}(0-)=\ell^{(r)}_{0}]$
$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[(1-e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}})1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+\alpha^{(r)}_{e}(e^{r\theta(\ell^{(r)}_{0}+1)}-1)\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-1)1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$=-\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\qquad+\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.44)
Define ${\mathcal{E}}^{(r)}_{1}$ is defined as
${\mathcal{E}}^{(r)}_{1}$ $=2^{-1}\mathbb{E}^{(r)}_{e}[\{(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}\}1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.45)
Applying (4.8) and
$e^{r\theta}e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}=(1+r\theta+o(r\theta))\{1-r\theta(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})$
$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}]+o((r\theta)^{2})\}$
$\quad=1-r\theta(-1+\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-2^{-1}(r\theta)^{2}[((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}]$
$\qquad+S_{3}(\widehat{R}^{(r)}_{e})o((r\theta)^{2}),$
to (4.3) and recalling the definition of $\Delta^{(r)}_{1}$, we have
$D^{(r)}(\theta,0)=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(-r\theta\Delta^{(r)}_{1}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{1})+o((r\theta)^{2}),$ (4.46)
because $\mathbb{E}^{(r)}_{s}[S_{3}(\widehat{R}^{(r)}_{e})1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
Similarly, define ${\mathcal{E}}^{(r)}_{2}$ as
${\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{e}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (4.47)
Then, $D^{(r)}(0,\theta)$ is computed from (4.2) as
@@
$\quad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}[(e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}-1)1(L^{(r)}(0-)=\ell^{(r)}_{0})]+\alpha^{(r)}_{e}(1-e^{r\theta(\ell^{(r)}_{0}+1)})\mathbb{P}^{(r)}_{s}[L^{(r)}(0)=\ell^{(r)}_{0}]$
@@
$=\alpha^{(r)}_{e}e^{r\theta\ell^{(r)}_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s}}1(L^{(r)}(0-)=\ell^{(r)}_{0})]-e^{r\theta}\mathbb{E}^{(r)}_{s}[e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}}1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}$
$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}\mathbb{E}^{(r)}_{e}[1+r\theta\mu^{(r)}\widehat{R}^{(r)}_{s-}-2^{-1}(r\theta)^{2}[(\mu^{(r)}\sigma^{(r)}_{s})^{2}(\mu^{(r)}\widehat{R}^{(r)}_{s-})-(\mu^{(r)}\widehat{R}^{(r)}_{s-})^{2}]1(L^{(r)}(0-)=\ell^{(r)}_{0})]$
$\quad-(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}[1-r\theta(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})$
$\qquad-2^{-1}(r\theta)^{2}[(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}]1(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}+o((r\theta)^{2})$
$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(r\theta\Delta^{(r)}_{2}+(r\theta)^{2}{\mathcal{E}}^{(r)}_{2})+o((r\theta)^{2}).$ (4.48)
(Step 3)   In this step, we derive asymptotic expansions for $\psi^{(r)}_{i}(\theta_{i})$ for $i=1,2$. First, note that (4.35) and (4.36) can be written as
$\alpha^{(r)}_{e}\Delta^{(r)}_{1}=-rb_{1}\mu\varphi^{(r)}_{1}(0)+c^{(r)}_{1}\mu^{(r)}\mathbb{P}[L^{(r)}=0]+o(r),$ (4.49)
$\alpha^{(r)}_{e}\Delta^{(r)}_{2}=r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r).$ (4.50)
Then, substituting (4.46) into the asymptotic BAR (4.3) with $\theta_{2}=0$ and applying (4.49), we have
$\frac{1}{2}rc_{1}\mu\sigma_{1}^{2}\big{(}\beta_{1}-\theta_{1}\big{)}\psi^{(r)}_{1}(\theta_{1})=c_{1}\mu\mathbb{P}[L^{(r)}=0]-e^{\theta_{1}\ell_{0}}\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-r\theta_{1}{\mathcal{E}}^{(r)}_{1})+\theta_{1}o(r),$
$\quad=c_{1}\mu(1-e^{\theta_{1}\ell_{0}})\mathbb{P}[L^{(r)}=0]+re^{\theta_{1}\ell_{0}}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\theta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+\theta_{1}o(r),$ (4.51)
Similarly, substituting (4.3) into (4.3) with $\theta_{1}=0$ and applying (4.50), we have
$\frac{1}{2}rc_{2}\mu\sigma_{2}^{2}\big{(}\beta_{2}-\theta_{2}\big{)}\psi^{(r)}_{2}(\theta_{2})=\alpha^{(r)}_{e}e^{\theta_{2}\ell_{0}}\big{(}\Delta^{(r)}_{2}+r\theta_{2}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r)$
$\quad=re^{\theta_{2}\ell_{0}}\big{(}\mu b_{2}\varphi^{(r)}_{2}(0)+\theta_{2}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{2}\big{)}+\theta_{2}o(r).$ (4.52)
We next compute $\mathbb{P}[L^{(r)}=0]$ for (4.3).
@@
$\mathbb{P}[L^{(r)}=0]=\begin{cases}\frac{r}{2c_{1}\mu\ell_{0}}(c_{1}\mu\sigma_{1}^{2}\varphi^{(r)}_{1}(0)+2\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1})+o(r),&b=0,\\ \frac{re^{\beta_{1}\ell_{0}}}{c_{1}\mu(e^{\beta_{1}\ell_{0}}-1)}\big{(}b_{1}\mu\varphi^{(r)}_{1}(0)+\beta_{1}\alpha^{(r)}_{e}{\mathcal{E}}^{(r)}_{1}\big{)}+o(r),&b\not=0.\end{cases}$ (4.53)